# SCOI: Syntax-augmented Coverage-based In-context Example Selection for Machine Translation

### v1

@@ Abstract
In-context learning (ICL) greatly improves the performance of large language models (LLMs) on various down-stream tasks, where the improvement highly depends on the quality of demonstrations. In this work, we introduce syntactic knowledge to select better in-context examples for machine translation (MT). We propose a new strategy, namely Syntax-augmented COverage-based In-context example selection (SCOI), leveraging the deep syntactic structure beyond conventional word matching. Specifically, we measure the set-level syntactic coverage by computing the coverage of polynomial terms with the help of a simplified tree-to-polynomial algorithm, and lexical coverage using word overlap. Furthermore, we devise an alternate selection approach to combine both coverage measures, taking advantage of syntactic and lexical information. We conduct experiments with two multi-lingual LLMs on six translation directions. Empirical results show that our proposed SCOI obtains the highest average COMET score among all learning-free methods, indicating that combining syntactic and lexical coverage successfully helps to select better in-context examples for MT.
@@ 1 Introduction
Our code will be publicly available after the publication of this paper.
@@ 3.3 Measure of Set-level Lexical Coverage
3.4 Combined Syntactic and Lexical Coverage
@@ 4.3 Large Language Models
4.4 Prompt Template
The number of in-context examples is set to 4 in our experiments. We use the same prompt template used in Kumar et al. (2023) for XGLM and that in He et al. (2024) for Alpaca. Please refer to Appendix D for details.
Other details of our implementation can be found in Appendix E.
@@ 5.1 Main Results
Table 2: COMET scores of 4-shot ICL performance of SCOI and other methods for translation on all 6 directions of 2 language models. All the methods except CTQ Scorer are learning-free, which do not require task, language or LLM-specific training. "Avg." refers to the average score across all 6 directions. The highest scores among learning-free methods are in bold text.
Main results are shown in Table 2. SCOI obtains the highest COMET scores of 4 out of 6 translation directions and the highest average COMET score among all learning-free methods using both XGLM and Alpaca, showing competitive performance across language models. Using XGLM, SCOI outperforms the learning-based CTQ Scorer on "RU-EN", while using Alpaca, SCOI even outperforms CTQ Scorer on both "RU-EN" and "EN-RU". Note that Alpaca seems not good at generating Russian, which can be inferred from its poor random and BM25 baselines. But SCOI greatly improves its performance on "EN-RU" and shows amazing ability in teaching an LLM to better translate into a language that appeared less during training.
We observe that SCOI shows obvious preference across different languages. For example, it fails to benefit "FR-EN" but improves performance on "RU-EN". Besides different natures of different languages, this might be also due to different capabilities of syntax parsers for different languages. We find that when the parser performs poorly (e.g., the French parser), SCOI also performs less competitively, while a more powerful parser (e.g., the Russian one) leads to better performance of SCOI. Details of the relation between parser and SCOI’s performance can be found in Appendix F.
We also experiment on GPT-3.5 Ouyang et al. (2022), which is an API-based LLM. Results are presented in Appendix G.
@@ 5.3 Experiments with Different Selection Modes
We select the top-$k$ examples with the highest syntactic similarity based on the polynomial distance used in Liu et al. (2022) for each test input from the example database. Details of polynomial distance can be found in Appendix H.
@@
For the actual selection of $S_{\text{best}}$, we utilize the exact implementation of the greedy algorithm from Ye et al. (2023), originally proposed in Chen et al. (2018). Other details of DPP are presented in Appendix I.
@@ 5.5 Case Analysis
Table 6: An end-to-end German-to-English translation example. "Input & Gold" refers to the test input and the gold reference. "BM25 Prediction" refers to XGLM’s prediction given the test input and examples selected by BM25, which are shown in Appendix K. "Example-$i$" refers to the $i$-th example selected by SCOI. "SCOI Prediction" shows the predict of XGLM given the test input and the 4 in-context examples selected by SCOI.
@@
Table 6 also compares SCOI’s system output with that of BM25. BM25 fails to construct the proper syntactic structure when translating the German phrase "der Ausweitung des Handels in Asien" and turns it into a reported clause "that trade in Asia had been expanded", thus losing accuracy. Note that "der Ausweitung des Handels in Asien" (the expansion of trade in Asia) does not include temporal information and it could be a bygone, a current state or a future trend, while the result of BM25 assumes that it is something that happened in the past, which is inconsistent with the original meaning of the input sentence. However, SCOI, combining syntactic and lexical coverage, is able to output the exact noun phrase "the expansion of trade in Asia", which is consistent with the syntactic structure in the source German sentence and much more accurate in translation. For the complete end-to-end case of BM25, please refer to Appendix K.
@@ 6 Conclusion
In this work, we introduce syntactic information to in-context example selection for MT. We propose to measure set-level syntactic coverage with the set-level coverage of polynomial terms representing root-to-node paths in syntax trees based on a simplified tree-to-polynomial algorithm and select in-context example selection for MT using an alternate mode where syntactic and lexical coverage are considered alternately to combine information of syntax and word. Our proposed method obtains the highest average COMET score among all learning-free methods, indicating that combining syntactic and lexical coverage during in-context example selection is helpful for MT. We call on the NLP community to pay more attention to syntactic knowledge for syntax-rich tasks like MT.
@@ A.2 Our Simplified Algorithm
We then apply Equation 11 and 17 and ignore the constant factors to get
@@
we can continue the process recursively (in fact, each level of recursion corresponds to a layer in the tree) until each node has appeared on left-hand side and add together Equation 20, 22 and so on to get
@@ Appendix B BLEU Results
Table 9: BLEU scores of 4-shot ICL performance of SCOI and other methods for translation on all 6 directions of 2 language models. All the methods except CTQ Scorer are learning-free, which do not require task, language or LLM-specific training. "Avg." refers to the average score across all 6 directions. The highest scores among learning-free methods are in bold text.
@@ Appendix C The spaCy Models Used for Parsing
Appendix D Prompt Template
For XGLM, we use the same prompt template as used in Kumar et al. (2023):
    [source] sentence: [X_1]
    [target] sentence: [Y_1]
    [source] sentence: [X_k]
    [target] sentence: [Y_k]
    [source] sentence: [X]
    [target] sentence:
With the same symbols above, for Alpaca, we use the same prompt template as used in He et al. (2024):
Instruction: Translate the following
[source] text into [source].
[source]: [X_1]
[target]: [Y_1]
[source]: [X_k]
[target]: [Y_k]
[source]: [X]
[target]:
Appendix E Implementation Details
Appendix F Effect of Parser
@@
Appendix H Polynomial Distance
Note that we can write polynomial terms as term vectors as shown in Equation 5. In this way, a polynomial $P$ can be written as a set of term vectors $\mathcal{V}_{P}$. Then, following Liu et al. (2022), we compute the distance between two polynomials ($P$ and $Q$) as:
where ${\parallel s-t\parallel}_{1}$ is the Manhattan distance (Craw, 2017) between term vector $s$ and $t$.
Appendix I Details of DPPs
@@
Appendix J Effect of Measure of Coverage
Coverage Into EN Out of EN Avg.
DE FR RU DE FR RU
Cosine Similarity 64.35 71.54 53.89 45.41 55.36 46.06 56.10
Normalized Distance 64.67 71.26 54.08 44.87 55.31 46.47 56.11
Appendix K The Example of BM25

### v2

@@ Abstract
In-context learning (ICL) greatly improves the performance of large language models (LLMs) on various down-stream tasks, where the improvement highly depends on the quality of demonstrations. In this work, we introduce syntactic knowledge to select better in-context examples for machine translation (MT). We propose a new strategy, namely Syntax-augmented COverage-based In-context example selection (SCOI), leveraging the deep syntactic structure beyond conventional word matching. Specifically, we measure the set-level syntactic coverage by computing the coverage of polynomial terms with the help of a simplified tree-to-polynomial algorithm, and lexical coverage using word overlap. Furthermore, we devise an alternate selection approach to combine both coverage measures, taking advantage of syntactic and lexical information. We conduct experiments with two multi-lingual LLMs on six translation directions. Empirical results show that our proposed SCOI obtains the highest average COMET score among all learning-free methods, indicating that combining syntactic and lexical coverage successfully helps to select better in-context examples for MT. Our code is available at https://github.com/JamyDon/SCOI.
@@ 3.3 Measure of Set-level Lexical Coverage
3.4 Combining Syntactic and Lexical Coverage
@@ 4.3 Large Language Models
4.4 Implementation Details
The number of in-context examples is set to 4 in our experiments.
For XGLM, we use the same prompt template as used in Kumar et al. (2023):
    [source] sentence: [X_1]
    [target] sentence: [Y_1]
    [source] sentence: [X_k]
    [target] sentence: [Y_k]
    [source] sentence: [X]
    [target] sentence:
With the same symbols above, for Alpaca, we use the same template as used in He et al. (2024):
Instruction: Translate the following
[source] text into [target].
[source]: [X_1]
[target]: [Y_1]
[source]: [X_k]
[target]: [Y_k]
[source]: [X]
[target]:
@@ 5.1 Main Results
Table 2: COMET scores of 4-shot ICL performance of SCOI and other methods for translation on all 6 directions of 2 language models. The zero-shot baseline of each model is listed in the first row. All methods except CTQ Scorer are learning-free, which do not require task, language or LLM-specific training. "Avg." refers to the average score across all 6 directions. The highest scores among learning-free methods are in bold text.
Main results are shown in Table 2. SCOI obtains the highest COMET scores of 4 out of 6 translation directions and the highest average COMET score among all learning-free methods using both XGLM and Alpaca, showing competitive performance across language models. Using XGLM, SCOI outperforms the learning-based CTQ Scorer on "RU-EN", while using Alpaca, SCOI even outperforms CTQ Scorer on both "RU-EN" and "EN-RU". Note that Alpaca seems not good at generating Russian, and its performance gain with 4-shot random examples is fairly poor compared with the zero-shot baseline. But SCOI greatly improves its performance on "EN-RU" and shows amazing ability in teaching an LLM to better translate into a language that appeared less during training.
We observe that SCOI shows obvious preference across different languages. For example, it fails to benefit "FR-EN" but improves performance on "RU-EN". Besides different natures of different languages, this might be also due to different capabilities of syntax parsers for different languages. We find that when the parser performs poorly (e.g., the French parser), SCOI also performs less competitively, while a more powerful parser (e.g., the Russian one) leads to better performance of SCOI. Details of the relation between parser and SCOI’s performance can be found in Appendix D. On top of that, other factors like the nature of different languages’ syntax might also contribute to the fluctuations across languages, which we leave for future work.
XGLM’s zero-shot COMET scores on "out of English" directions are negative values. This might be due to that XGLM fails to follow the machine translation task in the prompt and sometimes produces a wrong language.
We also experiment on GPT-3.5 Ouyang et al. (2022), which is an API-based LLM. Results are presented in Appendix E.
@@ 5.3 Experiments with Different Selection Modes
We select the top-$k$ examples with the highest syntactic similarity based on the polynomial distance used in Liu et al. (2022) for each test input from the example database. Note that we can write polynomial terms as term vectors as shown in Equation 5. In this way, a polynomial $P$ can be written as a set of term vectors $\mathcal{V}_{P}$. Then, following Liu et al. (2022), we compute the distance between two polynomials ($P$ and $Q$) as:
where ${\parallel s-t\parallel}_{1}$ is the Manhattan distance (Craw, 2017) between term vector $s$ and $t$. Instances with top-$k$ lowest polynomial distances to the test input are used as the in-context examples.
@@
For the actual selection of $S_{\text{best}}$, we utilize the exact implementation of the greedy algorithm from Ye et al. (2023), originally proposed in Chen et al. (2018). Other details of DPP are presented in Appendix F.
@@ 5.4 Analysis on the Selection Order
5.5 Analysis on the Measure of Coverage
Coverage Into EN Out of EN Avg.
DE FR RU DE FR RU
Cosine Similarity 64.35 71.54 53.89 45.41 55.36 46.06 56.10
Normalized Distance 64.67 71.26 54.08 44.87 55.31 46.47 56.11
@@ 5.6 Case Analysis
Table 7: An end-to-end German-to-English translation example. "Input & Gold" refers to the test input and the gold reference. "BM25 Prediction" refers to XGLM’s prediction given the test input and examples selected by BM25, which are shown in Appendix G. "Example-$i$" refers to the $i$-th example selected by SCOI. "SCOI Prediction" shows the predict of XGLM given the test input and the 4 in-context examples selected by SCOI.
@@
Table 7 also compares SCOI’s system output with that of BM25. BM25 fails to construct the proper syntactic structure when translating the German phrase "der Ausweitung des Handels in Asien" and turns it into a reported clause "that trade in Asia had been expanded", thus losing accuracy. Note that "der Ausweitung des Handels in Asien" (the expansion of trade in Asia) does not include temporal information and it could be a bygone, a current state or a future trend, while the result of BM25 assumes that it is something that happened in the past, which is inconsistent with the original meaning of the input sentence. However, SCOI, combining syntactic and lexical coverage, is able to output the exact noun phrase "the expansion of trade in Asia", which is consistent with the syntactic structure in the source German sentence and much more accurate in translation. For the complete end-to-end case of BM25, please refer to Appendix G.
@@ 6 Conclusion
In this work, we introduce syntactic information to in-context example selection for MT. First, we measure set-level syntactic coverage with coverage of polynomial terms based on a simplified algorithm that converts syntax trees into polynomials. Then, we propose to select in-context examples for MT based on syntactic and lexical coverage alternately to combine information of syntax and word. Our proposed method obtains the highest average COMET score among all learning-free methods, indicating that combining syntactic and lexical coverage during in-context example selection is helpful for MT. We call on the NLP community to pay more attention to syntactic knowledge for syntax-rich tasks like MT.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62076008) and the Key Project of Natural Science Foundation of China (61936012).
@@ A.2 Our Simplified Algorithm
We then apply Equation 13 and 19 and ignore the constant factors to get
@@
we can continue the process recursively (in fact, each level of recursion corresponds to a layer in the tree) until each node has appeared on left-hand side and add together Equation 22, 24 and so on to get
@@ Appendix B BLEU Results
Table 10: BLEU scores of 4-shot ICL performance of SCOI and other methods for translation on all 6 directions of 2 language models. The zero-shot baseline of each model is listed in the first row. All the methods except CTQ Scorer are learning-free, which do not require task, language or LLM-specific training. "Avg." refers to the average score across all 6 directions. The highest scores among learning-free methods are in bold text.
@@ Appendix C The spaCy Models Used for Parsing
Appendix D Effect of Parser
@@
Appendix F Details of DPPs
@@
Appendix G The Example of BM25